# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v1

@@ Appendix A Appendix
Genre tags
We grouped genre tags into genre groups, as depicted in Table 5. Each row links a coarse “Group” label—such as Rock, Electronic, or Jazz/Blues—to the fine-grained “Genre tags” that appear in the metadata. These tags enumerate substyles (e.g., progressiverock, deephouse, acidjazz), which allows us to aggregate diverse representations inside each of the genre grops.
@@
Spectrogram examples
We visualize time–frequency structure in spectrograms to provide qualitative evidence of restoration behavior. Each figure shows the clean reference, the degraded input (e.g., reverberation-induced smearing or clipping distortion), and the output of SonicMaster. Figures 5, 6, 7, 8, and 9 compare clean, degraded, and restored spectrograms across selected scenarios (reverb, clipping, microphone transfer function, and clarity EQ).
@@
Prompts for each degradation type

### v2

@@ Appendix A Appendix
A.1 The Use of Large Language Models
We employed a Large Language Model to assist with reducing wordy paragraphs to help the paper fit in the page limit.
A.2 Genre tags
We grouped genre tags into genre groups, as depicted in Table 7. Each row links a coarse “Group” label—such as Rock, Electronic, or Jazz/Blues—to the fine-grained “Genre tags” that appear in the metadata. These tags enumerate substyles (e.g., progressiverock, deephouse, acidjazz), which allows us to aggregate diverse representations inside each of the genre groups.
@@ A.3 Degradation functions
A.4 Spectrogram examples
We visualize time–frequency structure in spectrograms to provide qualitative evidence of restoration behavior. Each figure shows the clean reference, the degraded input (e.g., reverberation-induced smearing or clipping distortion), and the output of SonicMaster. Figures 3, 4, 5, 6, 7, and 8 compare clean, degraded, and restored spectrograms across selected scenarios (reverb, clipping, microphone transfer function, and clarity EQ).
(a) Ground truth audio.
(b) Degraded by reverb (smearing).
(c) After SonicMaster: reverb removed.
Figure 4: Comparison of spectrograms: (a) ground truth, (b) degraded with reverb, and (c) the output of SonicMaster where smearing is removed. Prompt: “Please, reduce the strong echo in this song.”
@@ A.4 Spectrogram examples
[…]
Table 10: EQ Objective evaluation (average absolute error) – the lower, the better.
Model (MMDiT/DiT) Reverb Dynamics Amplitude Stereo
Small Big Mix Real Compressor Punch Clip Volume
[…]
Table 11: Objective evaluation: Reverb, Dynamics, Amplitude, and Stereo. Clip values are multiplied by 1000.
Model Single deg. Double+triple deg. All
[…]
Table 12: Objective evaluation: FAD, KL, SSIM, and PQ. For readability, KL values were multiplied by 1000.
A.5 Ablation on ODE Solvers
We evaluated Euler solvers with $1$, $10$ (baseline), and $100$ steps, plus a $10$-step 4th order Runge–Kutta (Dormand & Prince, 1980) solver. Tables A.4, A.4, and A.4 outline the results and highlight the trade-off across degradation categories. Euler-1 matches the baseline overall but is weaker on Boom, Microphone, Clip, all Reverb subtasks, and shows higher KL. Euler-100 boosts Reverb and Punch yet lowers every EQ score versus the 1-/10-step runs. Runge–Kutta-10 equals Euler-10 on most metrics and tops Clip, but its inference is significantly slower.
A.6 Prompts for each degradation type